Suppose that you want to estimate the relationship between peoples weight (W) and the number of times they eat out in a month (EO) where Ao is the intercept of the population regression line: β1 is the slope of the population regression line; ui įs the error term; and the subscript i runs over observations, i= 1, n For this, you collect data rom a random sample 01250 people. After anal zing he data, you determin na the c arance between peoples e and hen mber。 times they eat out in a month is 4.85 and the variance of the number of times people eat out in a month is 3.78. You also find that the mean weight of people in the sample is 61.58 kg and the mean number of times people eat out in a month is 2.25 The OLS estimator of the slope β1 is D (Round your answer to two decimal places.)
Answer #1

answer the OLS estimator of the slope \beta 1=1.2831

for the regression equation W=\beta0+\beta1*EO+ui

the OLS estimator of the slope \beta 1=Covaiance(W, EO)/Variance(EO)=4.85/3.78=1.2831

\beta0=mean(W)-\beta1*mean(EO)=61.58-1.2831*2.25=58.6930

following information has been given in the question

sample size=n=250, W=weight, EO=eat out

covariance(W, EO)=4.85

Variance(EO)=3.78

mean(W)=61.58

mean(EO)=2.25